# Initial-State Typicality in Quantum Relaxation

### v2

@@ Abstract
Introduction—Relaxation is a ubiquitous phenomenon in nature and a central process in open quantum systems. The relaxation time, also known as the mixing time, is a key figure of merit in diverse areas, including quantum information and computing [1, 2, 3], quantum simulations [4, 5, 6, 7], state preparation [8, 9, 10, 11], and quantum heat engines [12, 13, 14, 15, 16]. Over the past decades, influential lines of work employing rigorous mixing-time [17, 18, 19, 20, 21, 22, 23, 24, 25, 26] and spectral analysis [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] have clarified how the structure of the dynamics controls relaxation in open systems. These approaches have been instrumental in characterizing remarkable phenomena such as the cutoff phenomenon in classical and quantum Markov processes [17, 18], the Liouvillian skin effect [32, 33] and the rapid mixing [19, 20, 21, 22, 23, 24, 25]. A unifying aspect of these studies is their focus on identifying a single, initial-state-independent timescale to characterize a given dynamics. However, this approach overlooks the conceptually and practically important influence of the initial state on relaxation. Since every relaxation process begins from a specific state, its impact cannot be disregarded.
@@
where $\|A\|_{1}:={\rm tr}(\sqrt{A^{\dagger}A})$ is the trace norm. The additional constraint $\|R_{k}\|_{1}=1$ excludes any size-dependence from $R_{k}$ on the relaxation. We assume $\mathcal{L}$ has a unique stationary state $\rho_{\rm ss}$ (however, our analysis can be directly extended to cases where multi-stability exists). Then, given an arbitrary initial state $\rho_{0}$, the system density matrix at time $t$ is
@@
Example 1: Non-interacting dissipative $N$-qubit chain—We consider a non-interacting system consisting of $N$ damping qubits, with Hamiltonian $H=\sum_{i=1}^{N}E\sigma^{z}_{i}/2$ and local jump operators $J_{i,\downarrow}=\sqrt{\gamma_{1}}\,\sigma_{i}^{-},\quad J_{i,\uparrow}=\sqrt{\gamma_{0}}\;\sigma_{i}^{+}$. Here, $\sigma^{z}_{i}$ are Pauli operators of the $i$th particle, $\sigma_{i}^{-}=(|0\rangle\langle 1|)_{i}$ and $\sigma_{i}^{+}=(|1\rangle\langle 0|)_{i}$. The system can be analytically solved to illustrate the concentration phenomenon. In this case, it can be verified that [82] $\|L_{k}\|_{2}\sim\mathcal{O}(2^{N/2})\sim\mathcal{O}(\sqrt{d})$. Thus, the variances for both ensembles exponentially converge to 0 as ${\rm Var}_{U}(a_{2})\sim d^{-1}\sim 2^{-N},\ {\rm Var}(a_{2})\sim d^{-2}\sim 4^{-N}$. One can further add the dephasing noise $\sigma_{i}^{z}$, which will not change the scaling.
@@
Example 2: Dissipative transverse-field Ising model (TFIM)—We then numerically study an open TFIM of length $N$ with local dissipation. The Hamiltonian is $H=-J\sum_{i=1}^{N-1}\sigma_{i}^{z}\sigma_{i+1}^{z}-g\sum_{i=1}^{N}\sigma_{i}^{x}$, and jump operators are the same as Example 1 (with a physical constraint $\gamma_{1}/\gamma_{0}=e^{\beta E},$ where $\beta=1/k_{B}T$ is the inverse temperature of multiple identical heat baths independently coupled to each qubit). Notably, $\rho_{\rm ss}$ may not be an equilibrium state in this example. We find that the concentration of $a_{2}$ is robust in this model, across different parameter regimes ($J,\ g,\ \beta$), the details of a specific case is shown in End Matter. The robustness is reminiscent of the fact that there is no thermal phase transition in one-dimensional spin systems and there may be a connection to be explored.
@@
Sufficiently rapid relaxation may imply mode concentration—Rapid mixing, a well-known phenomenon states that the relaxation timescale for a system with size $n$ is of the order $\mathcal{O}(\log[{\rm poly}(n)])$, can imply the concentration of $a_{2}$. We prove that $\|L_{k}\|_{2}\leq 2\sqrt{d}\max_{\rho_{0}}|a_{k}|$ for any $k$. This inequality yields upper bounds for the variance: ${\rm Var}_{U}(a_{k})\leq 4(\max_{\rho_{0}}|a_{k}|)^{2}d^{-1}$ and ${\rm Var}(a_{k})\leq 4(\max_{\rho_{0}}|a_{k}|)^{2}d^{-2}$. Indeed, rapid mixing implies $\max_{\rho_{0}}|a_{2}|$ is at most of the order $\mathcal{O}({\rm poly}(\log d))$ under an assumption used in [34], ensuring the variances vanish as $d\to\infty$. The rapid mixing has been proven for diverse systems [19, 20, 21, 22, 24, 87]. Interestingly, rapid mixing implies stability of relaxation [19], and concentration also implies more stable relaxation modes, which indicates deep connections between them.
@@
However, fast relaxation is not a necessary condition for typicality. Since the closing of Liouvillian gap is sufficient for a slow relaxation, the slow relaxation alone cannot imply $O_{k}$ grows fast. Notably, a recent work provided an example for the typicality in slow relaxation [88] with a different normalization $\|R_{k}\|_{2}=1$. In their case, $\|L_{2}\|_{2}\approx\|\mathbb{I}_{d}\|=\sqrt{d}$ and thus ${\rm Var}(a_{2})\sim d^{-1}$.
@@
where $p_{\max/\min}(\rho_{\beta})$ are maximum/minimum eigenvalues of $\rho_{\beta}$ and $\Delta E:=E_{\rm max}-E_{\rm min}$ is the energy range of $H_{0}$. Here, by QDBC we refer to the Davies map [89, 90] or the Kubo–Martin–Schwinger (KMS) condition [22] with respect to $\rho_{\beta}$  [^3] See [82] for details of these conditions. It has been shown that, if a Lindbladian satisfies these conditions, $\rho_{\beta}=e^{-\beta H_{0}}/{\rm tr}(e^{-\beta H_{0}})$ is its steady state. Note that $H_{0}$ may not coincide with the system Hamiltonian $H$ of the Lindbladian, there could be a shift term, e.g., the Lamb shift [24].. Lindbladians satisfying the KMS condition have been constructed [8, 92] and prove critical to quantum simulations [7].
@@
Universality of the typicality for random mixed states—Beyond specific examples, we here provide general arguments, using insight from random matrix theory. For a large class of non-Hermitian random matrix such as complex Ginibre ensemble and random matrix whose entries are independent and identically distributed, the typical scaling of $O_{k}$ for bulk eigenvalues and large $d$ was rigorously established as $O_{k}\sim\mathcal{O}(d)$ [94, 95]. For random Lindbladian, this scaling has also been confirmed [96, 97]. This cannot guarantee the universality of the typicality for pure initial states.
However, if the initial state is a random mixed state, we expect that at least the concentration of the dominant relaxation mode ($a_{2}$) would be a universal behavior. The reason is that $\lambda_{2}$ should be classified as the edge eigenvalue. Indeed, for many random matrix models, the typical scaling of $O_{k}$ of an edge eigenvalue $\lambda_{k}$ was shown to be $\mathcal{O}(\sqrt{d})$ [98, 99]. Although for random Lindbladian, currently there is no rigorous proof, a typical scaling slower than $d$ can still be anticipated, which at least assures the concentration of $a_{2}$ for random mixed states. It is known that edge eigenvalues have larger eigenvalue spacing, which makes them more stable (smaller condition number). In a recent paper [100], it was shown that a larger eigenvalue spacing implies a smaller dimensional scaling of the condition number $O_{k}$.
@@
Re-evaluation of the efficiency of quantum Gibbs samplers—The mixing time is essential in evaluating the efficiency of a quantum Gibbs sampler [8, 92, 7] for efficient state preparations and quantum simulations. In addition to the conventional (worst-case) metric, it is natural to also report the typical relaxation benchmark—namely the typical relaxation time defined above. With initial-state typicality, establishing efficiency may amount to showing typical rapid mixing, i.e., that the typical relaxation time scales favorably, which is a weaker and more representative requirement than conventional rapid mixing. Even if a sampler is initialized atypically, a light randomization can push it into the typical set, as discussed above.
@@
Our work opens several promising avenues for future research. A concrete next step is to construct explicit physical models that exhibit a sharp transition into the typicality regime. Further questions concern the fluctuation-dominated regime where typicality fails: are there other universal statistical laws governing relaxation, perhaps revealed in the distribution of relaxation times? The potential universality of the typicality itself, hinted at by random matrix theory, also warrants deeper investigation. Additionally, the concentration of slow modes implies that associated metastable states [111] also exhibit typicality, potentially impacting quantum computation that leverages metastability [112, 113]. Perhaps most importantly, the typical relaxation time provides a powerful new tool. It may enable proofs of rapid mixing under broader conditions and redefine performance benchmarks across quantum science, from state preparation to quantum simulations.

### v3

@@ Abstract
Introduction—Relaxation is a ubiquitous phenomenon in nature and a central process in open quantum systems. The relaxation time, also known as the mixing time, is a key figure of merit in diverse areas, including quantum information and computing [1, 2, 3], quantum simulations [4, 5, 6, 7], state preparation [8, 9, 10, 11, 12], and quantum heat engines [13, 14, 15, 16, 17]. Over the past decades, influential lines of work employing rigorous mixing-time [18, 19, 20, 21, 22, 23, 24, 25, 26, 27] and spectral analysis [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40] have clarified how the structure of the dynamics controls relaxation in open systems. These approaches have been instrumental in characterizing remarkable phenomena such as the cutoff phenomenon in classical and quantum Markov processes [18, 19], the Liouvillian skin effect [33, 34] and the rapid mixing [20, 21, 22, 23, 24, 25, 26]. A unifying aspect of these studies is their focus on identifying a single, initial-state-independent timescale to characterize a given dynamics. However, this approach overlooks the conceptually and practically important influence of the initial state on relaxation. Since every relaxation process begins from a specific state, its impact cannot be disregarded.
@@
where $\|A\|_{1}:={\rm tr}(\sqrt{A^{\dagger}A})$ is the trace norm. The additional constraint $\|R_{k}\|_{1}=1$ excludes any size-dependence from $R_{k}$ on the relaxation. We assume $\mathcal{L}$ has a unique stationary state $\rho_{\rm ss}$ (however, our analysis can be directly extended to cases where multistability exists [82]). Then, given an arbitrary initial state $\rho_{0}$, the system density matrix at time $t$ is
@@
Example 1: Non-interacting dissipative $N$-qubit chain—We consider a non-interacting system consisting of $N$ damped qubits, with Hamiltonian $H=\sum_{i=1}^{N}E\sigma^{z}_{i}/2$ and local jump operators $J_{i,\downarrow}=\sqrt{\gamma_{1}}\,\sigma_{i}^{-},\quad J_{i,\uparrow}=\sqrt{\gamma_{0}}\;\sigma_{i}^{+}$. Here, $\sigma^{z}_{i}$ are Pauli operators of the $i$th particle, $\sigma_{i}^{-}=(|0\rangle\langle 1|)_{i}$ and $\sigma_{i}^{+}=(|1\rangle\langle 0|)_{i}$. The system can be analytically solved to illustrate the concentration phenomenon. In this case, it can be verified that [84] $\|L_{k}\|_{2}\sim\mathcal{O}(2^{N/2})\sim\mathcal{O}(\sqrt{d})$. Thus, the variances for both ensembles exponentially converge to 0 as ${\rm Var}_{U}(a_{2})\sim d^{-1}\sim 2^{-N},\ {\rm Var}(a_{2})\sim d^{-2}\sim 4^{-N}$. One can further add the dephasing noise $\sigma_{i}^{z}$, which will not change the scaling.
@@
Example 2: Dissipative transverse-field Ising model (TFIM)—We then numerically study an open TFIM of length $N$ with local dissipation. The Hamiltonian is $H=-J\sum_{i=1}^{N-1}\sigma_{i}^{z}\sigma_{i+1}^{z}-g\sum_{i=1}^{N}\sigma_{i}^{x}$, and jump operators are the same as Example 1 (with a physical constraint $\gamma_{1}/\gamma_{0}=e^{\beta E},$ where $\beta=1/k_{B}T$ is the inverse temperature of multiple identical heat baths independently coupled to each qubit). Notably, $\rho_{\rm ss}$ may not be an equilibrium state in this example. The steady state is unique in both this model and Example 1, as the jump operators satisfy the sufficient condition for uniqueness [65]. We find that the concentration of $a_{2}$ is robust in this model, across different parameter regimes ($J,\ g,\ \beta$), the details of a specific case are shown in End Matter. The robustness is reminiscent of the fact that there is no thermal phase transition in one-dimensional spin systems and there may be a connection to be explored.
@@
Sufficiently rapid relaxation may imply mode concentration—Rapid mixing, a well-known phenomenon stating that the relaxation timescale for a system with size $n$ is of the order $\mathcal{O}(\log n)$, can imply the concentration of $a_{2}$. We prove that $\|L_{k}\|_{2}\leq 2\sqrt{d}\max_{\rho_{0}}|a_{k}|$ for any $k$. This inequality yields upper bounds for the variance: ${\rm Var}_{U}(a_{k})\leq 4(\max_{\rho_{0}}|a_{k}|)^{2}d^{-1}$ and ${\rm Var}(a_{k})\leq 4(\max_{\rho_{0}}|a_{k}|)^{2}d^{-2}$. Indeed, rapid mixing implies $\max_{\rho_{0}}|a_{2}|$ is at most of the order $\mathcal{O}({\rm poly}(\log d))$ under an assumption used in [35], ensuring the variances vanish as $d\to\infty$. The rapid mixing has been proven for diverse systems [20, 21, 22, 23, 25, 89]. Interestingly, rapid mixing implies stability of relaxation [20], and concentration also implies more stable relaxation modes, which indicates deep connections between them.
@@
However, fast relaxation is not a necessary condition for typicality. Since the closing of Liouvillian gap is sufficient for a slow relaxation, the slow relaxation alone cannot imply $O_{k}$ grows fast.
@@
where $p_{\max/\min}(\rho_{\beta})$ are maximum/minimum eigenvalues of $\rho_{\beta}$ and $\Delta E:=E_{\rm max}-E_{\rm min}$ is the energy range of $H_{0}$. Here, by QDBC we refer to the Davies map [90, 91] or the Kubo–Martin–Schwinger (KMS) condition [23] with respect to $\rho_{\beta}$  [^3] See [84] for details of these conditions. It has been shown that, if a Lindbladian satisfies these conditions, $\rho_{\beta}=e^{-\beta H_{0}}/{\rm tr}(e^{-\beta H_{0}})$ is its steady state. Note that $H_{0}$ may not coincide with the system Hamiltonian $H$ of the Lindbladian, there could be a shift term, e.g., the Lamb shift [25].. Lindbladians satisfying the KMS condition have been constructed [8, 93] and have proven critical to quantum simulations [7]. Notably, we assume that $\rho_{\beta}$ is the unique steady state, which is natural in the context of thermalization, although the upper bound remains valid in the presence of multistability.
@@
Finally, our framework offers a representative benchmark for assessing the efficiency of quantum Gibbs samplers [8, 93, 7], complementing conventional worst-case metrics (see End Matter).
@@
Our work opens several promising avenues for future research. A concrete next step is to construct explicit physical models that exhibit a sharp transition into the typicality regime. The potential universality of the typicality itself, hinted at by random matrix theory, also warrants deeper investigation (see End Matter). Additionally, the concentration of slow modes implies that associated metastable states [105] also exhibit typicality, potentially impacting quantum computation that leverages metastability [106, 107]. It is also interesting to extend our framework to broader settings. For instance, adapting this framework to bosonic systems with infinite-dimensional Hilbert spaces is not yet straightforward and would likely require a reformulation of the key ingredients (e.g., random ensembles) via physically motivated truncations [108, 109, 110]. Perhaps most importantly, whether the typical relaxation time can help relax some of the conditions required for proving rapid mixing deserves further investigation, as it may provide alternative performance benchmarks in applications ranging from state preparation to quantum simulation  [^5] Specifically, a large separation between typical-case and worst-case mixing is expected when the maximal overlap with the slowest mode is large while the typical overlap is suppressed, a scenario exemplified by systems exhibiting the TSME. If the TSME can be established for specific low-temperature thermalization (e.g., in one-dimensional spin systems as suggested by the numerical results in Example 2), it may uncover concrete examples of typical rapid mixing even where conventional rapid mixing fails..
@@
Data availability statement—The code used to generate the data in Fig. 1 is open available at  [^6] https://github.com/Ruicheng-12/Initial-state-typicality-in-quantum-relaxation.